2. Suppose the consumer consume good t&S and get the utility 𝑠 1 4 𝑡 3 4 and the consumer spent his total income birr 300 to purchase good t at price = 10 birr &S at price=20 birr per unit. Then based on the above information answer the following questions. A. Calculate the utility maximizing level of consumption of good t&S? (1 point) B. Calculate marginal utility of t at utility maximizing level of consumption of good t&S level and interpret it? (2 points) C. Calculate MRSst at utility maximizing level of consumption of good t&S level and interpret it? (2 points) D. Show the result in graph? (1 point


1
Expert's answer
2022-03-09T11:03:06-0500

a)


"\frac {MU_t}{P_t}=\frac {34}{10}"


"=3.4"


b)


"\frac {MU_s}{P_s}=\frac {14}{20}"


"=0.7"


c)

Hence, the consumer is not at the optimal position at 10 units of t and 20 units of s.


d)